Two weights w 1 and w 2 are suspended from the ends of a light string passing over a smooth fixed pulley.If the pulley is pulled up at an acceleration g,the tension in the string will be how much???

Two weights w1 and w2 are suspended from the ends of a light string passing over a smooth fixed pulley.If the pulley is pulled up at an acceleration g,the tension in the string will be how much??? 

Grade:12th pass

1 Answers

Arun
5 years ago
Dear Sucheta
 
g`=2gso w1`=2w1 and w2`=2w2Using T=2m1.m2.g`/(m1+m2) and substituting m1=w1/g and m2=w2/g and g` =2gWe get, tension = 4w1w2/(w1+w2)
 
Regards
Arun